what does this question mean find three consecutive numbers such that three times the middle number is 12 more than the sum of the other two numbers[hint.let the three nubers be x,x+1,etc.]
the topic is solving problems with eqations

Answer by ewatrrr(24785)

three nubers be x,x+1, x+2
3(x+1) = (x + (x+2))+12
x = 11
three consecutive numbers are: 11, 12, 13
And...
36 = 24 +12